OpenStudy (anonymous):

MEDAL!! True or False: The graph of y = sin x is increasing on the interval pi/2 , 3pi/2 . Explain your answer

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

find the derivative

ganeshie8 (ganeshie8):

if the derivative is POSITIVE in given interval, then the function is INCREASING

ganeshie8 (ganeshie8):

y = sin(x) dy/dx = ?

ganeshie8 (ganeshie8):

familiar with derivatives ha ?

ganeshie8 (ganeshie8):

Is this question related to Calculus ?

OpenStudy (anonymous):

so it is true then

OpenStudy (anonymous):

its pre-calc

ganeshie8 (ganeshie8):

it is False actually

OpenStudy (anonymous):

how would I explain that one lol

OpenStudy (anonymous):

Oh ok I see that

ganeshie8 (ganeshie8):

you might explain it like below : the graph of y = sin(x) has a highest value of 1 at x = pi/2 and it is continually decreasing upto x = 3pi/2; thus the function y = sin(x) is decreasing in the interval (pi/2, 3pi/2)

OpenStudy (anonymous):

Thanks! I just have 1 left on my study guide

ganeshie8 (ganeshie8):

so the given statement is FALSE< the graph is NOT INCREASING,

ganeshie8 (ganeshie8):

good luck !